OpenStudy (anonymous):

The Ratio of the Surface Area in square meters to volume cubic meters is 1:5. Find the radius of the sphere.

OpenStudy (anonymous):

\(Area=4 \pi r^2, Volume=\frac{4}{3} \pi r^3\) you are given the ratio of s.a : v = 1:5 solve: \(1:5 = S.A : V\\1:5=4 \pi r^2:\frac{4}{3} \pi r^3\\5(4 \pi r^2)=\frac{4}{3} \pi r^3\\20 \pi r^2=\frac{4}{3} \pi r^3\\60\pi r^2=4\pi r^3\) divide both sides by \(\pi r^2\) \(60=4r\) divide both sides by 4 and you'll get it..

OpenStudy (anonymous):

Refer to the attachement.

OpenStudy (anonymous):

thanks